# Overcoming Saturation in Density Ratio Estimation by Iterated Regularization

### v1

@@ 6.2 Methods
We test our suggested iterated regularization strategy for the four different baseline approaches in Example 1: KuLSIF (Kanamori et al., 2009), LR (Bickel et al., 2009), Exp (Menon and Ong, 2016) and SQ (Menon and Ong, 2016). Except for KuLSIF, which has a closed form solution, we rely on the CG method implemented in Python Scipy (Virtanen et al., 2020) with Polak-Ribiere line search scheme (Hager and Zhang, 2006).
@@
To get the full picture of the impact on different deep domain adaptation methods, we compute ensemble model candidates for $11$ domain adaptation algorithms: Adversarial Spectral Kernel Matching (AdvSKM) (Liu and Xue, 2021), Deep Domain Confusion (DDC) (Tzeng et al., 2014), Correlation Alignment via Deep Neural Networks (Deep-Coral) (Sun et al., 2017), Central Moment Discrepancy (CMD) (Zellinger et al., 2017), Higher-order Moment Matching (HoMM) (Chen et al., 2020), Minimum Discrepancy Estimation for Deep Domain Adaptation (MMDA) (Rahman et al., 2020), Deep Subdomain Adaptation (DSAN) (Zhu et al., 2021), Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016), Conditional Adversarial Domain Adaptation (CDAN) (Long et al., 2018), A DIRT-T Approach to Unsupervised Domain Adaptation (DIRT) (Shu et al., 2018) and Convolutional deep Domain Adaptation model for Time-Series data (CoDATS) (Wilson et al., 2020).
@@ 7 Conclusion and Future Work
One important future aspect is parameter choice, especially under the requirement of reliable methods with theoretical error guarantees, since our approach introduces the number of iterations as a new parameter.
@@ Appendix A Proof of Theorem 1
We start with the following common technical assumption, essentially requiring the reproducing kernel Hilbert space $\mathcal{H}$ to be rich enough and regular in a weak sense (cf. Caponnetto and De Vito [2007], Marteau-Ferey et al. [2019b], Gizewski et al. [2022], Zellinger et al. [2023]):
@@
To obtain fast error rates, we rely on a general self-concordance property of loss functions [Marteau-Ferey et al., 2019b], which follows from Assumption 3. To this end, we denote by $A[k,h]$ the evaluation of an operator $A:\mathcal{H}\times\mathcal{H}\to\mathbb{R}$ at the arguments (functions) $k,h\in\mathcal{H}$.
Assumption 5 (generalized self-concordance [Marteau-Ferey et al., 2019b]).
@@
Under Assumption 5, an upper bound on the excess risk of iteratively regularized kernel regression can be provided [Beugnot et al., 2021], which reads in our notation as follows.
Lemma 2 (Beugnot et al. [2021, Theorem 1, first part]).
@@
We refer to the supplementary material of Beugnot et al. [2021] for the exact values of $L_{0},N,C_{\mathrm{var}},C_{\mathrm{bias}}$, which depend on $r,\alpha,S,R,t,\delta$ and the distribution $\rho$ but not on the sample size $m+n$.
@@ Appendix B Optimization Error
This Section is aimed to provide a more detailed sketch on how one can obtain the error bound Eq. (3). The first quantity we need to consider in this context is the so called Newton decrement, as developed in Marteau-Ferey et al. [2019a]. Starting from $f^{\lambda,0}_{{\bf z}}=\bar{f}^{\lambda,0}_{{\bf z}}=0$, we define the following expressions for $k>0$ :
$f^{\lambda,t+1}_{{\bf z}}$ $=\operatorname*{arg\,min}_{f\in\mathcal{H}}\frac{1}{|{\bf z}|}\sum_{i=1}^{m+n}\ell(y_{i},f(x_{i}))+\frac{\lambda}{2}\left\lVert f-f^{\lambda,t}_{{\bf z}}\right\rVert^{2}:=\operatorname*{arg\,min}_{f\in\mathcal{H}}\widehat{\mathcal{R}}_{\lambda}^{t}(f)$ (14)
$\nu_{\lambda}^{t+1}(f)$ $:=\left\|{\bf H}_{\lambda}(f)^{-\frac{1}{2}}\nabla\widehat{\mathcal{R}}_{\lambda}^{t}(f)\right\|$ (15)
$\bar{f}^{\lambda,t+1}_{{\bf z}}$ $\approx\operatorname*{arg\,min}_{f\in\mathcal{H}}\frac{1}{|{\bf z}|}\sum_{i=1}^{m+n}\ell(y_{i},f(x_{i}))+\frac{\lambda}{2}\left\lVert f-\bar{f}^{\lambda,t}_{{\bf z}}\right\rVert^{2}:=\operatorname*{arg\,min}_{f\in\mathcal{H}}\bar{\mathcal{R}}_{\lambda}^{t}(f)$ (16)
$\bar{\nu}_{\lambda}^{t+1}(f)$ $:=\left\|{\bf H}_{\lambda}(f)^{-\frac{1}{2}}\nabla\bar{\mathcal{R}}_{\lambda}^{t}(f)\right\|$ (17)
for a numerical approximation $\bar{f}_{{\bf z}}^{\lambda,t}$ of $f^{\lambda,t}_{{\bf z}}$ computed as described in Subsection 4.1. Next we need to enforce a bound on the true Newton decrement in Eq. (15) when we only have access to ¯$\bar{\mathcal{R}}_{\lambda}^{t}$. This has been done in Beugnot et al. [2021, Proposition 1], and we repeat their result in our notation:
@@
[Beugnot et al., 2021, Proposition 1] Let $\varepsilon>0$ the target precision. Assume that we can solve each sub-problem with precision $\bar{\epsilon}_{k}$ :
@@
To compute this estimator in practice, one would need to solve $t$ optimization problems with decreasing precision. As the CG-algorithm that we are applying (cf. Section 4.1) has error complexity $\log(\varepsilon)$ (see Remark (3)), the complexity of applying this algorithm iteratively would only be $t$ times bigger than computing a single iteration, so achieving the required small magnitudes of the weighted gradient norms Eq. (17)–(15) is computationally feasible. Analyzing the computational burden in terms of the number of samples required to compute the numerical solution of Eq. (4) would be an interesting avenue for future work. To conclude our investigations on how the theoretical guarantees change by the use of inexact solvers, we can adapt Beugnot et al. [2021, Proposition 2] with the same arguments as used in our proof of Theorem  1:
@@ Appendix C Derivation of Iterated KuLSIF
as in Kanamori et al. [2012a] (following the representer Theorem), plugging this into Eq. (18) and taking the derivatives w.r.t. $\alpha_{j}^{t+1}$ and $\beta_{l}^{t+1}$ and setting them to 00, we get as analytic expressions for the coefficients:
@@ Appendix D Dataset with Known Density Ratios
To study the accuracy of the iteratively regularized estimations of density ratios, we follow investigations of Kanamori et al. [2012a] and Nguyen et al. [2010a] to construct high dimensional data with exact known density ratios. Thereby, we both extend and generalize existing settings. For example datasets such as Ringnorm or Twonorm Breiman [1996] are special cases of a Gaussian Mixture approach with mean and covariance parameters of the respective mixture modes set accordingly. Similarly, experiments in Nguyen et al. [2010a] can be modeled by a Gaussian Mixture distribution. We make the task more complex by randomly sampling the number, the weights, the expectations and the covariances of mixture components of the underlying distributions from a space with higher dimension ($50$) compared to existing work. More specifically, we sample the expectations from $[0,0.5]^{50}$ and the weights from $[0,1]$ with additional normalization. Each distribution (source, target) gets assigned to a different mixture model. In each experiment there is a maximum component count of $4$ components. The randomly sampled $n\in\{1,2,3\}$ is the number of components for the source and $4-n$ for the target distribution.
@@ E.1 Known density ratios
For each Gaussian mixture dataset we draw $5000$ samples from the underlying distributions. The selection and evaluation of the compared density ratio estimation method is done in a typical train/val/test split approach with split ratios $64/16/20$ respectively. The regularization (hyper) parameter $\lambda$ is selected from $\{10^{-6},10^{-5},\ldots,10^{4}\}$ and for each experiment $10$ replicates are carried out. Accordingly, the number of iteration steps is selected from $\{1,2,\ldots,10\}$ based on the respective loss metric. We follow Kanamori et al. [2012a] in using the Gaussian kernel with kernel width set according to the median heuristic Schölkopf and Smola [2002] for all compared density ratio estimation methods.
@@ E.2 Domain Adaptation
The computation of the results for the domain adaptation benchmark experiment is based on gradient-based training of overall 9174 models where we built parts of our results on the codebase of Dinu et al. [2023] which results in the following specifics.
@@
Following Dinu et al. [2023] we used $11$ domain adaptation methods from the AdaTime [Ragab et al., 2023] benchmark. For each of these methods and domain adaptation modalities (text, image, sensory) we evaluate all $8$ density ratio estimation on all $22$ domain adaptation scenarios. For our experiments we follow the same model implementation and experimental setup as in Dinu et al. [2023] which results in the usage of fully connected networks for Amazon Reviews and a pretrained ResNet-18 backbone for MiniDomainNet. For training and selecting the density ratio estimation methods within this pipeline we perform an additional train/val split of $80/20$ on the datasets that are used for training the domain adaption methods. The regularization (hyper) parameter $\lambda$ is selected from $\{10^{-6},10^{-5},\ldots,10^{4}\}$ and due to higher computational effort the number of iterations $t$ is selected from $\{1,5,10\}$ based on the respective Bregman objective (1). We follow Kanamori et al. [2012a] in using the Gaussian kernel with kernel width set according to the median heuristic [Schölkopf and Smola, 2002] for all compared density ratio estimation methods. Each experiment is run $3$ times. To test the performance of the compared methods the classification accuracy on the respective test sets of the target distribution is evaluated. We use the ensemble approach introduced in Dinu et al. [2023] as discussed in Section 6.
E.3 Matrix Inversion
As shown in Dinu et al. [2023], the solution of minimization problem (9) involves the inversion of a matrix that can be ill-conditioned such that similar models can cause numerical instabilities due to limited compute precision. We similarly rely on computing the pseudo-inverse with the Python Package numpy.linalg.pinv. It involves a singular-value based regularization scheme based on a threshold rcond where we average our results based on rcond values from $10^{-4}$ to $10^{-1}$ to separate numerical inversion problems and performance comparisons of density ratio estimation methods.

### v2

@@ 6.2 Methods
We test our suggested iterated regularization strategy for the four different baseline approaches in Example 1: KuLSIF (Kanamori et al., 2009), LR (Bickel et al., 2009), Exp (Menon & Ong, 2016) and SQ (Menon & Ong, 2016). Except for KuLSIF, which has a closed form solution, we rely on the CG method implemented in Python Scipy (Virtanen et al., 2020) with Polak-Ribiere line search scheme (Hager & Zhang, 2006).
@@
To get the full picture of the impact on different deep domain adaptation methods, we compute ensemble model candidates for $11$ domain adaptation algorithms: Adversarial Spectral Kernel Matching (AdvSKM) (Liu & Xue, 2021), Deep Domain Confusion (DDC) (Tzeng et al., 2014), Correlation Alignment via Deep Neural Networks (Deep-Coral) (Sun et al., 2017), Central Moment Discrepancy (CMD) (Zellinger et al., 2017), Higher-order Moment Matching (HoMM) (Chen et al., 2020), Minimum Discrepancy Estimation for Deep Domain Adaptation (MMDA) (Rahman et al., 2020), Deep Subdomain Adaptation (DSAN) (Zhu et al., 2021), Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016), Conditional Adversarial Domain Adaptation (CDAN) (Long et al., 2018), A DIRT-T Approach to Unsupervised Domain Adaptation (DIRT) (Shu et al., 2018) and Convolutional deep Domain Adaptation model for Time-Series data (CoDATS) (Wilson et al., 2020).
@@ 7 Conclusion and Future Work
One important future aspect is parameter choice, especially under the requirement of reliable methods with theoretical error guarantees, since our approach introduces the number of iterations as a new parameter. We also plan to improve our (time-invariant) time-series domain adaptation using the time-dependent Bregman divergence approach of Kim et al. (2023).
A reference implementation of the methods presented in this paper is available at: https://github.com/lugruber/dre_iter_reg.
Acknowledgements
We thank Otmar Scherzer, Sergei Pereverzyev, Stefan Kindermann, Marius-Constantin Dinu and four anonymous reviewers for helpful comments. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proof of Theorem 1
We start with the following common technical assumption, essentially requiring the reproducing kernel Hilbert space $\mathcal{H}$ to be rich enough and regular in a weak sense (cf. Caponnetto & De Vito (2007); Marteau-Ferey et al. (2019b); Gizewski et al. (2022); Zellinger et al. (2023)):
@@
To obtain fast error rates, we rely on a general self-concordance property of loss functions (Marteau-Ferey et al., 2019b), which follows from Assumption 3. To this end, we denote by $A[k,h]$ the evaluation of an operator $A:\mathcal{H}\times\mathcal{H}\to\mathbb{R}$ at the arguments (functions) $k,h\in\mathcal{H}$.
Assumption 5 (generalized self-concordance (Marteau-Ferey et al., 2019b)).
@@
Under Assumption 5, an upper bound on the excess risk of iteratively regularized kernel regression can be provided (Beugnot et al., 2021), which reads in our notation as follows.
Lemma 2 (Beugnot et al. (2021, Theorem 1, first part)).
@@
We refer to the supplementary material of Beugnot et al. (2021) for the exact values of $L_{0},N,C_{\mathrm{var}},C_{\mathrm{bias}}$, which depend on $r,\alpha,S,R,t,\delta$ and the distribution $\rho$ but not on the sample size $m+n$.
@@ Appendix B Optimization Error
This Section is aimed to provide a more detailed sketch on how one can obtain the error bound Eq. (3). The first quantity we need to consider in this context is the so called Newton decrement, as developed in Marteau-Ferey et al. (2019a). Starting from $f^{\lambda,0}_{{\bf z}}=\bar{f}^{\lambda,0}_{{\bf z}}=0$, we define the following expressions for $k>0$ :
$f^{\lambda,t+1}_{{\bf z}}$ $=\operatorname*{arg\,min}_{f\in\mathcal{H}}\frac{1}{|{\bf z}|}\sum_{i=1}^{m+n}\ell(y_{i},f(x_{i}))+\frac{\lambda}{2}\left\lVert f-f^{\lambda,t}_{{\bf z}}\right\rVert^{2}:=\operatorname*{arg\,min}_{f\in\mathcal{H}}\widehat{\mathcal{R}}_{\lambda}^{t}(f),\quad\nu_{\lambda}^{t+1}(f):=\left\|{\bf H}_{\lambda}(f)^{-\frac{1}{2}}\nabla\widehat{\mathcal{R}}_{\lambda}^{t}(f)\right\|,$ (14)
$\bar{f}^{\lambda,t+1}_{{\bf z}}$ $\approx\operatorname*{arg\,min}_{f\in\mathcal{H}}\frac{1}{|{\bf z}|}\sum_{i=1}^{m+n}\ell(y_{i},f(x_{i}))+\frac{\lambda}{2}\left\lVert f-\bar{f}^{\lambda,t}_{{\bf z}}\right\rVert^{2}:=\operatorname*{arg\,min}_{f\in\mathcal{H}}\bar{\mathcal{R}}_{\lambda}^{t}(f),\quad\bar{\nu}_{\lambda}^{t+1}(f):=\left\|{\bf H}_{\lambda}(f)^{-\frac{1}{2}}\nabla\bar{\mathcal{R}}_{\lambda}^{t}(f)\right\|,$ (15)
for a numerical approximation $\bar{f}_{{\bf z}}^{\lambda,t}$ of $f^{\lambda,t}_{{\bf z}}$ computed as described in Subsection 4.1. Next we need to enforce a bound on the true Newton decrement in Eq. (14) when we only have access to ¯$\bar{\mathcal{R}}_{\lambda}^{t}$. This has been done in Beugnot et al. (2021, Proposition 1), and we repeat their result in our notation:
@@
(Beugnot et al., 2021, Proposition 1) Let $\varepsilon>0$ the target precision. Assume that we can solve each sub-problem with precision $\bar{\epsilon}_{k}$ :
@@
To compute this estimator in practice, one would need to solve $t$ optimization problems with decreasing precision. As the CG-algorithm that we are applying (cf. Section 4.1) has error complexity $\log(\varepsilon)$ (see Remark (3)), the complexity of applying this algorithm iteratively would only be $t$ times bigger than computing a single iteration, so achieving the required small magnitudes of the weighted gradient norms Eq. (15)–(14) is computationally feasible. Analyzing the computational burden in terms of the number of samples required to compute the numerical solution of Eq. (4) would be an interesting avenue for future work. To conclude our investigations on how the theoretical guarantees change by the use of inexact solvers, we can adapt Beugnot et al. (2021, Proposition 2) with the same arguments as used in our proof of Theorem  1:
@@ Appendix C Derivation of Iterated KuLSIF
as in Kanamori et al. (2012a) (following the representer Theorem), plugging this into Eq. (16) and taking the derivatives w.r.t. $\alpha_{j}^{t+1}$ and $\beta_{l}^{t+1}$ and setting them to 00, we get as analytic expressions for the coefficients:
@@ Appendix D Dataset with Known Density Ratios
To study the accuracy of the iteratively regularized estimations of density ratios, we follow investigations of Kanamori et al. (2012a) and Nguyen et al. (2010a) to construct high dimensional data with exact known density ratios. Thereby, we both extend and generalize existing settings. For example datasets such as Ringnorm or Twonorm (Breiman, 1996) are special cases of a Gaussian Mixture approach with mean and covariance parameters of the respective mixture modes set accordingly. Similarly, experiments in Nguyen et al. (2010a) can be modeled by a Gaussian Mixture distribution. We make the task more complex by randomly sampling the number, the weights, the expectations and the covariances of mixture components of the underlying distributions from a space with higher dimension ($50$) compared to existing work. More specifically, we sample the expectations from $[0,0.5]^{50}$ and the weights from $[0,1]$ with additional normalization. Each distribution (source, target) gets assigned to a different mixture model. In each experiment there is a maximum component count of $4$ components. The randomly sampled $n\in\{1,2,3\}$ is the number of components for the source and $4-n$ for the target distribution.
@@ E.1 Known density ratios
For each Gaussian mixture dataset we draw $5000$ samples from the underlying distributions. The selection and evaluation of the compared density ratio estimation method is done in a typical train/val/test split approach with split ratios $64/16/20$ respectively. The regularization (hyper) parameter $\lambda$ is selected from $\{10^{-6},10^{-5},\ldots,10^{4}\}$ and for each experiment $10$ replicates are carried out. Accordingly, the number of iteration steps is selected from $\{1,2,\ldots,10\}$ based on the respective loss metric. We follow (Kanamori et al., 2012a) in using the Gaussian kernel with kernel width set according to the median heuristic (Schölkopf & Smola, 2002) for all compared density ratio estimation methods.
@@ E.2 Domain Adaptation
The computation of the results for the domain adaptation benchmark experiment is based on gradient-based training of overall 9174 models where we built parts of our results on the codebase of Dinu et al. (2023) which results in the following specifics.
@@
Following Dinu et al. (2023) we used $11$ domain adaptation methods from the AdaTime (Ragab et al., 2023) benchmark. For each of these methods and domain adaptation modalities (text, image, sensory) we evaluate all $8$ density ratio estimation on all $22$ domain adaptation scenarios. For our experiments we follow the same model implementation and experimental setup as in Dinu et al. (2023) which results in the usage of fully connected networks for Amazon Reviews and a pretrained ResNet-18 backbone for MiniDomainNet. For training and selecting the density ratio estimation methods within this pipeline we perform an additional train/val split of $80/20$ on the datasets that are used for training the domain adaption methods. The regularization (hyper) parameter $\lambda$ is selected from $\{10^{-6},10^{-5},\ldots,10^{4}\}$ and due to higher computational effort the number of iterations $t$ is selected from $\{1,5,10\}$ based on the respective Bregman objective (1). We follow Kanamori et al. (2012a) in using the Gaussian kernel with kernel width set according to the median heuristic (Schölkopf & Smola, 2002) for all compared density ratio estimation methods. Each experiment is run $3$ times. To test the performance of the compared methods the classification accuracy on the respective test sets of the target distribution is evaluated. We use the ensemble approach introduced in Dinu et al. (2023) as discussed in Section 6.
E.3 Further Ablation Studies
SOTA Domain Adaptation
We compare our method to the state-of-the-art results reported in Dinu et al. (2023). Table 3 shows the results for each domain adaptation method (rows) averaged over 12 domain adaptation tasks and three seeds for the Amazon Reviews dataset. While KuLSIF without our proposed iteration only achieves $0.787$ and does not outperform (Dinu et al., 2023), our proposed iterated KuLSIF improves (Dinu et al., 2023) from $0.788$ to $0.790$.
Domain Adaptation: Amazon Reviews
DA-Method Dinu et al. (2023) KuLSIF Iterated KuLSIF (ours)
HoMM $0.788(\pm 0.010)$ $0.777(\pm 0.009)$ $0.777(\pm 0.010)$
AdvSKM $0.780(\pm 0.009)$ $0.779(\pm 0.011)$ $0.780(\pm 0.008)$
DIRT $0.787(\pm 0.008)$ $0.787(\pm 0.011)$ $0.794(\pm 0.011)$
DDC $0.780(\pm 0.010)$ $0.780(\pm 0.010)$ $0.780(\pm 0.010)$
CMD $0.794(\pm 0.009)$ $0.790(\pm 0.010)$ $0.794(\pm 0.008)$
MMDA $0.787(\pm 0.011)$ $0.786(\pm 0.011)$ $0.789(\pm 0.010)$
CoDATS $0.796(\pm 0.009)$ $0.795(\pm 0.012)$ $0.798(\pm 0.011)$
Deep-Coral $0.785(\pm 0.009)$ $0.784(\pm 0.009)$ $0.784(\pm 0.009)$
CDAN $0.788(\pm 0.010)$ $0.787(\pm 0.010)$ $0.790(\pm 0.010)$
DANN $0.797(\pm 0.009)$ $0.794(\pm 0.013)$ $0.800(\pm 0.009)$
DSAN $0.795(\pm 0.009)$ $0.794(\pm 0.011)$ $0.800(\pm 0.009)$
Avg. $0.788(\pm 0.009)$ $0.787(\pm 0.011)$ $0.790(\pm 0.010)$
Table 3: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews over three different random initialization of model weights and 12 domain adaptation tasks.
Extension to Deep Learning
We follow Section 4.1 of Rhodes et al. (2020) to analyze the sample efficiency for challenging (different) densities. More precisely, we sample from an extremely peaked Gaussian $p\sim\mathcal{N}(0,10^{-6})$ and a broad Gaussian $q\sim\mathcal{N}(0,1)$. The density ratio estimators are based on both standard logistic regression and multi-layer networks. In Figure 3 (left) we can see that our iteration method improves sample efficiency for logistic regression. Additionally, it can be seen that our approach can be extended to neural network based models for which sample efficiency is improved as well.
Figure 3: Sample efficiency curves for various density ratio estimators and approaches, the error is measured by L1-norm. Left: Comparison logistic regression and multi-layer network logistic regression density ratio estimators both without and with iteration. Right: Comparison of logistic regression with its iterated version, the telescoping (tel) approach from Rhodes et al. (2020) and the telescoping approach combined with our iterative method.
Combining Iteration with Telescoping
We extend the previous experiment such that we sample from the same respective distributions and combine the telescoping framework of Rhodes et al. (2020) with our iteration method. In Figure 3 (right) it can be seen that telescoping improves the non-iterated density ratio estimator as in Rhodes et al. (2020). Additionally, the combination of telescoping and our iteration method further improves sample efficiency of both approaches.